The next 9 questions are related to the titration of 25.00 mL of a 0.1000 M acetic acid solution with 0.1000 M KOH.

1.What is the initial pH of the analyte solution?

2.What volume of KOH is required to reach the equivalence point of the titration (in mL)?

3.How many mmol of the salt are present at the equivalence point? (ANALYTICAL AMOUNT, NOT EQUILIBRIUM AMOUNT)

4.What is the volume of the solution at the equivalence point (in mL)?

5.What is the molar concentration of the salt at the equivalence point? (ANALYTICAL CONCENTRATION, NOT EQUILIBRIUM CONCENTRATION)

6.What is the pOH at the equivalence point?

7.What is the pH at the equivalence point?

8.How many mL of KOH have been added at the half-equivalence point?

9.What is the pH at the half-equivalence point?
Answer #1

Please go through detailed solution provided.... Images attached-Titration of 25mL, O. IM CH₂COOH (weak acid) by 0.1M KOH (strong Base ) CHCOOH + KOH CH₃Cook + H2O 8-1. Initial solution 0.1+ Ho 7-3. CHCOOH + KOH Chlook Salt mmoles of salt = mmoles of the cool = Mx V (mL) Mamolarity = 0.1 95 mmol - 2.5 mmol 8-4 nipH = 5 - 0 2 5) + (-2 + 7) = 7+ 4.75 + Ź (-113) = 7 + 2:37 - 0.65 8.72 [PA + POH_14 pH = 14- 8.72 = 5.28 8-7 - pH _ 8.72) cal